The bank balance sheet below lists the categories of assets and liabilities, along with the total amount of each
category, and the amount in each category that is "interest rate sensitive" or repriced within one year.
Calculate the existing Dollar Gap for the bank. Next, calculate the effect (change) on this bank's Net Interest
Income if interest rates fall or decrease by 1 percentage point or 100 bp. "%" denotes either the current
interest rate earned earned or paid on the designated asset or liability category. RSA and RSL denote interest
rate sensitive assets and liabilities, respectively, that are repriced within one year. Note, you do not have to
include either the existing or new level of Net Interest Income, as I'm only asking for the change in Net Interest
Income. Please show your work in order to receive credit.

  

Assets

Amount

%

RSA

Liabilities and Equity

Amount

%

RSL

Cash

80

0.0%

Non-interest deposits

100

0.0%

Securities

250

6.0%

100

NOW checking

200

2.0%

Loans, net

620

7.0%

248

MMDA

300

4.0%

300

Fed funds sold

0

CDs

150

5.0%

75

Non-earning assets

50

Fed Funds purchased

150

3.0%

150

$1,000

Equity

100

$1,000
Answer #1

Dollar GAP = RSA-RSL RSA: 100 +248 = 348 RS2 = 300 +75+ 150=525 Dollar GAP=RSA-RSL = 348-525 :-177 present Nil= (250 x 6%) +